# When does bottom-up beat top-down in hierarchical community detection?

### v2

@@ 4.1 Chernoff-Hellinger Divergence
The hardness of separating nodes that belong to a primitive community $a\in\mathcal{L}_{\mathcal{T}}$ from nodes in community $b\in\mathcal{L}_{\mathcal{T}}$ is quantified by the Chernoff-Hellinger divergence, denoted by $\mathrm{CH}(a,b)=\mathrm{CH}(a,b,\pi,p,\mathcal{T})$ and defined by444 AS15a originally defined the Chernoff-Hellinger divergence as $\mathrm{CH}_{AS}(a,b)\ =\ \sup_{t\in(0,1)}(1-t)\sum_{c\in\mathcal{L}_{\mathcal{T}}}\pi_{c}\left(tp_{ac}+(1-t)p_{bc}-p_{ac}^{t}p_{bc}^{1-t}\right).$ When $p_{ab}=o(1)$, we note that $(1-t)\operatorname{D}_{t}\left(\operatorname{Ber}\left(p_{ac}\right)\|\operatorname{Ber}\left(p_{bc}\right)\right)\ =\ (1+o(1))\left(tp_{ac}+(1-t)p_{bc}-p_{ac}^{t}p_{bc}^{1-t}\right)$, hence our definition of $\mathrm{CH}(a,b)$ in (4.1) coincides with the original definition $\mathrm{CH}_{AS}(a,b)$ (up to second-order terms).
@@
The condition for achieving exact recovery of the communities displays a phase transition. More precisely, when $p_{ab}=\Theta(\log N/N)$, exact recovery of the primitive communities is possible if $\lim\frac{NI}{\log N}>1$ and is impossible when $\lim\frac{NI}{\log N}<1$ AS15a ; AS15b .
@@
Consider a BTSBM with $K=2^{d}$ communities and $\pi_{a}=1/K$ for all $a\in[K]$. Suppose that for all $t\in[d]\colon p_{t}=a_{t}\log N/N$, with $a_{t}$ positive constants. Simple computations (see for example (Lei, 20, Lemma 6.6)) yield that
@@ 4.2 Exact recovery at Intermediate Depths
Although the exact recovery of the primitive communities is feasible when the quantity $I$, as defined in Equation (4.2), exceeds the threshold of $\log N/N$, this insight alone does not provide the requirements for achieving exact recovery of the super-communities of higher depth within the hierarchy. In this section, we answer this question by determining the precise information-theoretic threshold for achieving the exact recovery of super-communities at a specific depth. Furthermore, we demonstrate that a bottom-up algorithm successfully attains this threshold.
We say that a node $u$ of $\mathcal{T}$ is at depth $q$ if its distance from the root is $q$, or if $u$ is a leaf [^5] The second condition ($u$ is a leaf) is only needed if the binary tree is not perfect, as we highlight in Example 3.. We denote by $\mathcal{S}_{q}$ (resp., $\mathcal{S}_{\leq q}$) the set of nodes at depth $q$ (resp., at a depth less than or equal to $q$), i.e.,
$\mathcal{S}_{q}\ =\ \Big{\{}u\in\mathcal{T}\colon\big{(}|u|=q\big{)}\text{ or }\big{(}|u|\leq q\text{ and }u\text{ is a leaf}\big{)}\Big{\}}\quad\text{ and }\quad\mathcal{S}_{\leq q}\ =\ \bigcup\limits_{r\leq q}\mathcal{S}_{r}.$
@@
Because the tree of Figure 1(a) is not full and complete, the community $C_{10}$ is present at depth $2$ and at depth $3$. In contrast, for the tree of Figure 1(b), we have
@@
where the condition $\mathrm{lca}(a,b)\in\mathcal{S}_{\leq q-1}$ ensures that the lowest common ancestor of $a$ and $b$ has a depth less than or equal to $q-1$. When $q=d$, the minimum in Equation (4.5) is taken over all pairs of primitive communities, and the divergence $I_{d}$ is equal to the divergence $I$ defined in (4.2). Theorem 3 states the condition for recovering the communities at depth $q$.
@@
Let $G$ be an HSBM whose latent binary tree $\mathcal{T}$ has $K=2^{d}$ leaves. Let $q\in\{1,\cdots,d_{\mathcal{T}}\}$ and denote by $I_{q}$ the quantity defined in (4.5). The following holds:
@@
    exact recovery of the super-communities at depth $q$ is impossible if lim $\limsup\frac{NI_{q}}{\log N}<1$;
@@
    if lim $\liminf\frac{NI_{q}}{\log N}>1$ and $\min\limits_{u\in\mathcal{T}}p(u)=\omega(N^{-2})$, then the agnostic-degree-profiling algorithm of AS15b exactly recovers the super-communities at depth $q$.
Because $I_{q}$ is the minimum taken over all $\mathcal{S}_{\leq q-1}$, the quantity $I_{q}$ is non-decreasing in $q$. This is the reason why recovery at a higher intermediate depth $q$ is easier than recovery at a lower depth $q^{\prime}>q$. An extra condition $\min\limits_{u\in\mathcal{T}}p(u)=\omega(N^{-2})$ is present in point (ii) of Theorem 3 to rule out some particular cases [^6] For example, consider a toy BTSBM with 3 layers for which $p_{0}=0$, $p_{1}=N^{-10}$, $p_{2}=2\log N/N$ and $p_{3}=2.01\log N/N$. Although exact recovery at depth $3$ is impossible, exact recovery at depth $2$ is still possible. Nonetheless, exact recovery at depth 1 is impossible because, with a high probability, there are no edges between the super-communities having depth $2$..
We prove Theorem 3 in Appendix B.1. Although the quantity $I_{q}$ defined in (4.5) is generally hard to simplify, the following lemma provides a simple expression for BTSBMs.
@@
We prove Lemma 1 in Appendix B.2. In particular, Lemma 1 shows that when $q=d$, we have $I_{d}=K^{-1}\operatorname{D}_{1/2}\left(\operatorname{Ber}(p_{d-1}),\operatorname{Ber}(p_{d})\right)$, and hence we recover the divergence $I$ defined in (4.2) (see also Example 2).
@@ 5.1 Previous Work on Exact Recovery in HSBM
Early works by Dasgupta et al. DHKM (06) establish conditions for the exact recovery by top-down algorithms in an HSBM. Nonetheless, the recovery is ensured only in relatively dense regimes (specifically, average degree at least $\log^{6}N$), and the algorithm requires an unspecified choice of hyper-parameters. Balakrishnan et al. BXKS (11) show that a simple recursive spectral bi-partitioning algorithm recovers the hierarchical communities in a class of hierarchically structured weighted networks, where the weighted interactions are perturbed by sub-Gaussian noise. Nonetheless, the conditions require again a dense regime ((BXKS, 11, Assumption 1) states that all weights are strictly positive, hence the average degree scales as $\Theta(N)$), and the proposed algorithm has no stopping criterion. Recent work by Li et al. LLB+ (22) demonstrates that the same recursive spectral bi-partitioning algorithm exactly recovers the communities, when the average degree grows as $\Omega(\log^{2+\epsilon}N)$ for some $\epsilon>0$ (this condition can be further relaxed to $\Omega(\log N)$ by a refined analysis Lei (20)). The analysis of LLB+ (22) also allows for an unbounded number of communities and provides a consistent stopping criterion.
The linkage++ algorithm proposed by Cohen-Addad et al. CAKMT (17); CAKMTM (19) is a bottom-up algorithm that first estimates $K$ primitive communities by SVD and then successively merges them by using a linkage procedure. As SVD requires the number of clusters $K$ as an input, the authors propose to run the algorithms for every $K=2,\cdots,O(\log N)$ and to choose the optimal $K$ that leads to the hierarchical tree with the smallest Dasgupta-cost Das (16). Moreover, their analysis requires a dense setting in which the average degree grows as $\Omega(\sqrt{N\log N})$.
@@ 5.2 Top-Down HCD and Exact Recovery at Intermediate Depth
Sufficient conditions for the exact recovery of super-communities in BTSBMs by recursive spectral bi-partitioning algorithms were first established in LLB+ (22), under the assumptions that all connection probabilities have the same scaling factor $\delta_{N}$ (i.e., $p(t)=\Theta(\delta_{N})$ for all $t\in\mathcal{T}$) verifying $\delta_{N}=\Omega(\log^{2+\epsilon}N/N)$. The result was later refined in Lei (20) to remove the extra logarithmic factors and in LLL (21) to enable different scaling factors at different depths of the hierarchy. To the best of our knowledge, these are the only works involving recovery at intermediate depths of the hierarchy. To compare these results with Theorem 3, let us consider a BTSBM with $p(t)=a_{|t|}\log N/N$ where $a_{|t|}$ is constant independent of $N$. The average connection probability in a super-community of depth $q$ is then $\bar{a}_{q}\log N/N$ where $\bar{a}_{q}=\frac{1}{2^{d-q}}\left(a_{d}+\sum_{k=1}^{d-q}2^{k-1}a_{d-k}\right)$. Then (Lei, 20, Theorem 6.5) states that the recursive spectral bi-partitioning algorithm exactly recovers all the super-communities up to depth $\ell$ if
$\left(\sqrt{\bar{a}_{q}}-\sqrt{a_{q-1}}\right)^{2}>2^{q}\quad\text{ for all }q\in[\ell].$ (5.1)
@@
The condition $\left(\sqrt{\bar{a}_{q}}-\sqrt{a_{q-1}}\right)^{2}>2^{q}$ corresponds to the exact recovery threshold of a symmetric SBM with $2^{q}$ communities and intra-community (resp., inter-community) link probabilities $\bar{a}_{q}\log N/N$ (resp., $a_{q-1}\log N/N$). Although the depth $q$ of the hierarchy is composed of $2^{q}$ super-communities with average intra-connection probability $\bar{a}_{q}\log N/N$ and inter-connection probability $a_{q-1}\log N/N$, Condition (5.1) does not match the information-theoretic threshold derived in Theorem 3. Indeed, the structure of a super-community at depth $q$ is not simply an Erdős-Rényi random graph with connection probability $\bar{a}_{q}\log N/N$, but is instead an SBM composed of $2^{d-q}$ primitive communities, each of them interacting differently with the rest of the network.
@@
Let us compare (5.4) and (5.5). First, notice that for some choices of $a_{0},\cdots,a_{d}$, the function $q\mapsto J_{q}^{\mathrm{td}}$ given in (5.2) is not monotonically non-increasing [^7] Let $d=3$. For $(a_{0},a_{1},a_{2},a_{3})=(2.2,2.5,3,25)$ we have $(J_{1}^{\mathrm{td}},J_{2}^{\mathrm{td}},J_{3}^{\mathrm{td}})=(0.96,1.17,1.33)$, but for $(a_{0},a_{1},a_{2},a_{3})=(3,9,15,21)$ we have $(J_{1}^{\mathrm{td}},J_{2}^{\mathrm{td}},J_{3}^{\mathrm{td}})=(1.89,0.39,0.06)$. Finally, for $(a_{0},a_{1},a_{2},a_{3})=(2.2,2.4,4,22)$ we have $(J_{1}^{\mathrm{td}},J_{2}^{\mathrm{td}},J_{3}^{\mathrm{td}})=(0.85,1.02,0.90)$. Thus, the quantities $J_{1}^{\mathrm{td}},J_{2}^{\mathrm{td}},J_{3}^{\mathrm{td}}$ can be increasing, decreasing or interlacing., which contradicts the intuitive fact that the recovery at lower depths is harder than the recovery at higher depths. Moreover, although $J_{d}^{\mathrm{bu}}=J_{d}^{\mathrm{td}}$, Lemma 5 (deferred in Appendix B.3) shows that $J_{q}^{\mathrm{bu}}>J_{q}^{\mathrm{td}}$ for all $q\in[d-1]$, hence Condition (5.4) is strictly more restrictive than Condition (5.5). It is not clear whether the results of Lei (20) can be further improved as exact recovery by spectral bi-partitioning requires entry-wise concentration of the eigenvectors, which is established using sophisticated $\ell_{2\to\infty}$ perturbation bounds.

### v3

@@ 4.1 Chernoff-Hellinger Divergence
The hardness of separating nodes that belong to a primitive community $a\in\mathcal{L}_{\mathcal{T}}$ from nodes in community $b\in\mathcal{L}_{\mathcal{T}}$ is quantified by the Chernoff-Hellinger divergence,555abbe2015community originally defined the Chernoff-Hellinger divergence as $\mathrm{CH}_{AS}(a,b)\ =\ \sup_{t\in(0,1)}(1-t)\sum_{c\in\mathcal{L}_{\mathcal{T}}}\pi_{c}\left(tp_{ac}+(1-t)p_{bc}-p_{ac}^{t}p_{bc}^{1-t}\right).$ When $p_{ab}=o(1)$, we have $(1-t)\operatorname{D}_{t}\left(\operatorname{Ber}\left(p_{ac}\right)\|\operatorname{Ber}\left(p_{bc}\right)\right)=(1+o(1))\left(tp_{ac}+(1-t)p_{bc}-p_{ac}^{t}p_{bc}^{1-t}\right)$. Hence, our definition of $\mathrm{CH}(a,b)$ in (4.1) coincides with the original definition $\mathrm{CH}_{AS}(a,b)$ (up to second-order terms). denoted by $\mathrm{CH}(a,b)=\mathrm{CH}(a,b,\pi,p,\mathcal{T})$ and defined by
@@
The condition for achieving exact recovery of the communities displays a phase transition. More precisely, when $p_{ab}=\Theta(\log N/N)$, exact recovery of the primitive communities is possible if $\lim\frac{NI}{\log N}>1$ and is impossible when $\lim\frac{NI}{\log N}<1$ abbe2015community; abbe2015recovering.
@@
Consider a BTSBM with $K=2^{d}$ communities and $\pi_{a}=1/K$ for all $a\in[K]$. Suppose that for all $t\in[d]\colon p_{t}=a_{t}\log N/N$, with $a_{t}$ positive constants. Simple computations (see for example (lei2020unified, Lemma 6.6)) yield that
@@ 4.2 Exact recovery at Intermediate Depths
Although the exact recovery of the primitive communities is feasible when the quantity $I$, as defined in Equation (4.2), exceeds the threshold of $\log N/N$, this insight alone does not provide the requirements for achieving exact recovery of the super-communities at a smaller depth. In this section, we answer this question by determining the information-theoretic threshold for achieving the exact recovery of super-communities at a specific depth.
We say that a node $u$ of $\mathcal{T}$ is at depth $q$ if its distance from the root is $q$, or if $u$ is a leaf whose distance from the root is less than $q$. The second condition is only needed if the binary tree is not perfect, as we highlight in Example 3. We denote by $\mathcal{S}_{q}$ (resp., $\mathcal{S}_{\leq q}$) the set of nodes at depth $q$ (resp., at a depth less than or equal to $q$), i.e.,
[…]
@@
Because the tree of Figure 1(a) is not full and complete, the community $C_{10}$ is present at depth 22 and at depth $3$. In contrast, for the tree of Figure 1(b), we have
@@
where the condition $\mathrm{lca}(a,b)\in\mathcal{S}_{\leq q-1}$ ensures that the lowest common ancestor of $a$ and $b$ has a depth less than or equal to q1q-1. When $q=d$, the minimum in Equation (4.5) is taken over all pairs of primitive communities, and the divergence $I_{d}$ is equal to the divergence $I$ defined in (4.2). Theorem 3 states the condition for recovering the communities at depth $q$.
@@
Let $G$ be an HSBM and suppose Assumption 1 and 2 hold. Let $q\in\{1,\cdots,d_{\mathcal{T}}\}$ and denote by $I_{q}$ the quantity defined in (4.5). The following holds:
@@
    exact recovery of the super-communities at depth $q$ is impossible if $\limsup\frac{NI_{q}}{\log N}<1$;
@@
    if $\liminf\frac{NI_{q}}{\log N}>1$, then exact recovery of the super-communities at depth $q$ is possible.
Because $I_{q}$ is the minimum taken over all $\mathcal{S}_{\leq q-1}$, the quantity $I_{q}$ is non-increasing in $q$. This is the reason why exact recovery at a lower intermediate depth $q^{\prime}$ is easier than recovery at a higher depth $q>q^{\prime}$. We prove Theorem 3 in Appendix B.1. Although the quantity $I_{q}$ defined in (4.5) is generally hard to simplify, the following lemma provides a simple expression for BTSBMs.
@@
Lemma 1 shows that when $q=d$, we have $I_{d}=K^{-1}\operatorname{D}_{1/2}\left(\operatorname{Ber}(p_{d-1}),\operatorname{Ber}(p_{d})\right)$, and hence we recover the divergence $I$ defined in (4.2) (see also Example 2).
@@ 5.1 Previous Work on Exact Recovery in HSBM
Early works by Dasgupta et al. dasgupta2006spectral establish conditions for the exact recovery by top-down algorithms in an HSBM. Nonetheless, the recovery is ensured only in relatively dense regimes (specifically, average degree at least $\log^{6}N$), and the algorithm requires an unspecified choice of hyper-parameters. Balakrishnan et al. balakrishnan2011noise show that a simple recursive spectral bi-partitioning algorithm recovers the hierarchical communities in a class of hierarchically structured weighted networks, where the weighted interactions are perturbed by sub-Gaussian noise. Nonetheless, the conditions require again a dense regime ((balakrishnan2011noise, Assumption 1) states that all weights are strictly positive, hence the average degree scales as $\Theta(N)$), and the proposed algorithm has no stopping criterion. Recent work by Li et al. li2022hierarchical demonstrates that the same recursive spectral bi-partitioning algorithm exactly recovers the communities, when the average degree grows as $\Omega(\log^{2+\epsilon}N)$ for some $\epsilon>0$ (this condition can be further relaxed to $\Omega(\log N)$ by a refined analysis lei2020unified). The analysis of li2022hierarchical also allows for an unbounded number of communities and provides a consistent stopping criterion.
The linkage++ algorithm proposed by Cohen-Addad et al. cohen2017hierarchical; cohen2019hierarchical is a bottom-up algorithm that first estimates $K$ primitive communities by SVD and then successively merges them by using a linkage procedure. As SVD requires the number of clusters $K$ as an input, the authors propose to run the algorithms for every $K=2,\cdots,O(\log N)$ and to choose the optimal $K$ that leads to the hierarchical tree with the smallest Dasgupta-cost dasgupta2016cost. Moreover, their analysis requires a dense setting in which the average degree grows as $\Omega(\sqrt{N\log N})$.
Finally, in a recent paper, Fang and Rohe introduce the T-Stochastic Graph model fang2023t. This model captures hierarchical structures in networks by defining connection probabilities based on distances within an unrooted and potentially non-binary tree $T$. To reconstruct the latent hierarchy, the authors propose a bottom-up algorithm that integrates spectral clustering with the Neighbor-Joining algorithm. Unlike average linkage, which greedily merges the closest clusters, Neighbor-Joining iteratively selects pairs of clusters that minimize a criterion incorporating both pairwise distances and their average distances to all other nodes. Their key theoretical result, Theorem 4.3, establishes asymptotic recovery of the latent hierarchy when the average degree of the graph is $\Omega(\log^{11.1}n)$.
@@ 5.2 Top-Down HCD and Exact Recovery at Intermediate Depth
Exact recovery of communities at intermediate hierarchical depths is essential to the success of top-down algorithms, as current theoretical analyses li2022hierarchical; lei2020unified assume perfect bi-partitioning at each of the first $\ell$ recursive steps. This requires exact recovery of communities from depth 0 to depth $\ell$ in the hierarchy, which is only achievable above the information-theoretic threshold for exact recovery at depth $\ell$. Because exact recovery requires graph connectivity, the average degree must be $\Omega(\log N)$. In the following section, we show that the information-theoretic threshold is less stringent than the conditions in li2022hierarchical; lei2020unified. Although there may be room to improve on the conditions of li2022hierarchical; lei2020unified, current proof techniques cannot guarantee hierarchical recovery when the average degree is $o(\log N)$.
This raises the following question: Can we establish theoretical guarantees for top-down algorithms when the average degree grows slower than $\log N$? Current proof techniques are constrained by their requirement that all the $\ell$ first recursive steps must be error-free. This requirement ensures that the divisive algorithm recurse with HSBMs at each step. Indeed, consider an HSBM $G=(V,E)$ where a bi-partitioning algorithm produces partition $\widehat{\mathcal{C}}=(\widehat{C}_{1},\widehat{C}_{2})$ of $V$, yielding the subgraphs $\hat{G}_{1}=G[\widehat{C}_{1}]$ and $\hat{G}_{2}=G[\widehat{C}_{2}]$. When $\widehat{\mathcal{C}}$ exactly recovers depth-1 communities, we have either $\widehat{C}_{1}=C_{1}$ and $\widehat{C}_{2}=C_{1}$, or $\widehat{C}_{1}=C_{1}$ and $\widehat{C}_{2}=C_{2}$, ensuring that $\hat{G}_{1}$ and $\hat{G}_{2}$ are HSBMs.
This property no longer holds when $\widehat{\mathcal{C}}$ contains errors. Misclustered nodes create dependencies between edges in $\hat{G}_{1}$ (resp., in $\hat{G}_{2}$), as their assignments are determined by the bi-partitioning algorithm applied to $G$, so that $\hat{G}_{1}$ and $\hat{G}_{2}$ might bot be HSBMs any more. These correlations are complex to analyze, but we can bypass this difficulty and reframe the problem by viewing $\hat{G}_{1}$ and $\hat{G}_{2}$ as two perturbed HSBMs, where the perturbations result from connectivity modifications made by an adversary over a subset of nodes, whose size is upper-bounded by the number of misclustering errors. This formulation aligns with recent work on clustering robustness under adversarial perturbations cai2015robust; liu2022minimax; ding2023reaching, and could strengthen the theoretical guarantees for top-down algorithms. Furthermore, this adversarial framework explains why top-down algorithms using spectral clustering—a method known to be sensitive to arbitrary perturbations—are suboptimal.
5.3 Previous Work on Exact Recovery at Intermediate Depth
@@
Sufficient conditions for the exact recovery of super-communities in BTSBMs by recursive spectral bi-partitioning algorithms were established in li2022hierarchical, under the assumptions that all connection probabilities have the same scaling factor $\delta_{N}$ (i.e., $p(t)=\Theta(\delta_{N})$ for all $t\in\mathcal{T}$) verifying $\delta_{N}=\Omega(\log^{2+\epsilon}N/N)$. The result was later refined in lei2020unified to assume only that $\delta_{N}=\Omega(\log N/N)$ and in lei2021consistency to enable different scaling factors at different depths of the hierarchy. To compare these results with Theorem 3, let us consider a BTSBM with $p(t)=a_{|t|}\log N/N$ where $a_{|t|}$ is constant independent of $N$. The average connection probability in a super-community of depth $q$ is then $\bar{a}_{q}\log N/N$ where $\bar{a}_{q}=\frac{1}{2^{d-q}}\left(a_{d}+\sum_{k=1}^{d-q}2^{k-1}a_{d-k}\right)$. Then (lei2020unified, Theorem 6.5) states that the recursive spectral bi-partitioning algorithm exactly recovers all the super-communities up to depth $\ell$ if
[…]
@@
The condition $\left(\sqrt{\bar{a}_{q}}-\sqrt{a_{q-1}}\right)^{2}>2^{q}$ is the exact recovery threshold of a symmetric SBM with $2^{q}$ communities and intra-community (resp., inter-community) link probabilities $\bar{a}_{q}\log N/N$ (resp., $a_{q-1}\log N/N$). Although the depth $q$ of the hierarchy is composed of $2^{q}$ super-communities with average intra-connection probability $\bar{a}_{q}\log N/N$ and inter-connection probability $a_{q-1}\log N/N$, Condition (5.1) does not match the information-theoretic threshold derived in Theorem 3. Indeed, the structure of a super-community at depth $q$ is not an Erdős-Rényi random graph with connection probability $\bar{a}_{q}\log N/N$, but is instead an SBM with $2^{d-q}$ primitive communities.
@@
Let us compare (5.4) and (5.5). First, notice that for some choices of $a_{0},\cdots,a_{d}$, the function $q\mapsto J_{q}^{\mathrm{td}}$ given in (5.2) is not monotonically non-increasing, [^6] Let $d=3$. For $(a_{0},a_{1},a_{2},a_{3})=(2.2,2.5,3,25)$ we have $(J_{1}^{\mathrm{td}},J_{2}^{\mathrm{td}},J_{3}^{\mathrm{td}})=(0.96,1.17,1.33)$, but for $(a_{0},a_{1},a_{2},a_{3})=(3,9,15,21)$ we have $(J_{1}^{\mathrm{td}},J_{2}^{\mathrm{td}},J_{3}^{\mathrm{td}})=(1.89,0.39,0.06)$. Finally, for $(a_{0},a_{1},a_{2},a_{3})=(2.2,2.4,4,22)$ we have $(J_{1}^{\mathrm{td}},J_{2}^{\mathrm{td}},J_{3}^{\mathrm{td}})=(0.85,1.02,0.90)$. Thus, the quantities $J_{1}^{\mathrm{td}},J_{2}^{\mathrm{td}},J_{3}^{\mathrm{td}}$ can be increasing, decreasing or interlacing. which contradicts the intuitive fact that the recovery at larger depths is harder than the recovery at smaller depths. Moreover, although $J_{d}^{\mathrm{bu}}=J_{d}^{\mathrm{td}}$, Lemma 5 (deferred in Appendix B.3) shows that $J_{q}^{\mathrm{bu}}>J_{q}^{\mathrm{td}}$ for all $q\in[d-1]$, hence Condition (5.4) is strictly more restrictive than Condition (5.5). It is not clear whether the results of lei2020unified can be further improved as exact recovery by spectral bi-partitioning requires entry-wise concentration of the eigenvectors, which is established using sophisticated $\ell_{2\to\infty}$ perturbation bounds.